OpenStudy (anonymous):

Help me simplifiy it PLEASE!! Original question was: What is the simplified product of 3√10x times 2√15x? I need help on simplifying it!

OpenStudy (lgbasallote):

multiply 3 and 2 then multiply 10x and 15 x \[\Large (3)(2) \sqrt{(10x)(15x)}\]

OpenStudy (anonymous):

\[6\sqrt{150}\]

OpenStudy (anonymous):

\[6\sqrt{150x}\] 30√ 6x? or 30x √6